What mass of salt would have to be added to 1100.0 mL of water in order to raise the boiling point of the solution by 0.5 C? Assume the ideal van't Hoff factor for NaCl.
1
Expert's answer
2019-09-16T04:03:20-0400

dT = Kb * b * i

Kb for water = 0.513

b = 0.5/(0.513 * 2) = 0.48 (mol/kg)

b = n*i/m

n(NaCl) = b*m = 0.48 * 1.1 = 0.53 (mol)

m(NaCl) = n * M = 0.53 * 58.5 = 31 (g)


Answer: m(NaCl) = 31 g